OpenStudy (anonymous):

Line m has a slope of -1/5. What is the slope of a line that is perpendicular to line m?

OpenStudy (anonymous):

5 because they have to multiply to equal -1

OpenStudy (anonymous):

for a line to be perpendicular then you take the recipical of the slope : example if M=4 then perp M=1/4 and if m=1/6 then perp m=6

OpenStudy (anonymous):

it has to be the opposite recipricol

OpenStudy (anonymous):

so it would be -5?

OpenStudy (anonymous):

yes, sorry about that it is the neg

OpenStudy (anonymous):

yeah